OpenStudy (anonymous):

(x+4)^(2)=x^(2)+8 The solution set is?

OpenStudy (anonymous):

hey man

OpenStudy (anonymous):

you know how to find (x+4)^2

OpenStudy (anonymous):

(x+4)(x+4)

OpenStudy (anonymous):

(x+4)(x+4)

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

x^2+8x+16=x^2+8

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

8x+8=0

OpenStudy (anonymous):

8x=-8

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

I think it looking for two answers because it says seperate anser witha comma

OpenStudy (anonymous):

or should i just write the solution set is -1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well, unless you wrote the problem down wrong that should be it

OpenStudy (anonymous):

Thats the problem, all my other ones had two answers

OpenStudy (anonymous):

some problems will have 1, 2 , 3 ,4 answers ect..

OpenStudy (anonymous):

i can assure that this is the answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plug in -1

OpenStudy (anonymous):

in original prob